# A Survey on Trustworthiness in Foundation Models for Medical Image Analysis

### v1

@@ 4-B3 Fairness in Segmentation
As with base SAM itself, limitations are identified in the ability of MedSAM to perform reliably with underrepresented image modalities. In the context of MedSAM and medical image analysis, the concern comes from the fact that the training set largely consists of CT, MRI, or Endoscopy images. A similar issue is found in the variability of clinician needs in segmentation tasks; not all image modalities should be divided the same way. This presents a concern in the fairness of these segmentation models as they are used in rare clinical scenarios [9].
@@
MedSAM [9], a customization of the SAM fine-tuned on over 1M medical image-mask pairs covering 15 image modalities, has become a popular medical image analysis foundation model for segmentation. Xu et al. [39] examines the performance of MedSAM on a kidney CT scan set and looks for disparities in performance depending on patient characteristics. They use Dice similarity scores as their performance metric. No statistically significant relationship between patient age and performance is found, but a statistically significant negative correlation between patient BMI and performance has been noted. Xu et al. [39] also find that images from female patients yield better performance than from male patients.
4-C Report Generation
@@
4-C1 Privacy in Report Generation
Medical applications inherently come with significant concerns regarding patient data privacy. To satisfy privacy requirements, Thawkar et al. [11] de-identified the MIMIC-CXR dataset by removing patient information when training their model. To preserve the utility of the model trained on de-identified data, the authors used GPT-3.5-turbo to remove the de-defined symbols “__”, while preserving the original meaning of the reports.
4-C2 Robustness in Report Generation
@@
To improve the robustness of these models across different body regions, Zhong et al. [47] propose ChatRadio-Valuer, a model based on LLMs that learns generalizable representations from radiology reports of one institution and adapts to report generation tasks across different body regions (chest, abdomen, muscle-skeleton, head, maxillofacial, and neck). Another work on body-part robustness in report generation proposes a single-for-multiple (S4M) framework to make the model robust to six different body parts (chest, abdomen, knee, hip, wrist, and shoulder) while maintaining its performance or even outperforming other baseline methods [48]. This is achieved by incorporating general radiology knowledge with the radiology-informed knowledge aggregation branch and enhancing the cross-modal alignment by the implicit prior guidance branch.
Another robustness issue comes from the heterogeneity of different clinical institutions. To solve this problem, [47] starts with training on radiology reports from one institution to acquire knowledge of particular patterns and representations, and then it undergoes supervised fine-tuning using data from other numerous institutions. This approach ensures that the model can effectively handle a wide range of clinical scenarios and diagnostic tasks, improving its generalization ability and robustness to different institutions in real-world applications.
@@
In addition, medical images with blurry boundaries or noise can also lead the robustness concerns. To address the unstable generalization issues, an innovative SAM-guided dual-encoder architecture in MSMedCap is utilized to enable the capture of information with different granularities and overcome the unstable performance due to blurry boundaries, noise, and poor contrast in medical images [49]. Inspired by NEFTune [80], RO-LLaMA also enhances the model’s robustness and generalization abilities when faced with noisy inputs by adding random noise to the embedding vectors during training [61].
4-C3 Reliability in Report Generation
@@
Inheriting the issue of general foundation models, Vision Language Models (VLMs) based report generation systems also suffer from reliability issues. First of all, medical foundation models can make hallucinated references to non-existent prior reports when generating medical reports. One reason for this is that these models are trained on datasets of real-world patient reports that inherently refer to prior reports and lead to the incompatibility when solely inputting a medical image [54, 51]. To address this, [54] propose CXR-ReDonE to remove prior references from medical reports by using GPT-3 to rewrite the reports and using token classification to remove words referring to priors. Another study observe that when customizing general-purpose foundation models for medical report generation, they can hallucinate unintended text [50]. They attribute this to the use of cross-entropy loss for language modeling, which leads to the problem of exposure bias. One potential solution to address this issue is to leverage reinforcement learning with human (AI) feedback. In addition, Lee et al. [53] find that their proposed method for chest X-ray report generation may hallucinate reports with nonexistent findings. To mitigate this issue, they suggest some potential solutions, such as using higher quality/quantity training data, larger LLMs, or to strengthen the alignment between image and text modalities. Another work [55] proposes to enhance the reliability of radiology report generation by injecting additional knowledge with the current image in the generation process. Specifically, they fuse the weighted medical concept knowledge and reports for similar images with the current image features.
4-C4 Explainability in Report Generation
@@
Explainability is also a crucial concern in medical report generation. To improve the interpretability of report generation systems, Danu et al. [56] propose a two-step approach for generating the findings section of a radiology report from an automated interpretation of chest X-ray images: (1) detecting the abnormalities within the image using bounding-boxes with probabilities, and (2) harnessing the power of LLMs to translate the list of abnormalities into a Findings report. This two-step approach adds interpretability to the framework and aligns it with radiologists’ systematic reasoning during the review of CXRs.
@@
Using fine-grained vision-language models to provide alignment between image patches and texts which can match parts of the generated report with specific regions in the medical image can also improve explainability. However, using image patches with fixed sizes may result in incomplete representation of lesions which can occur in varying sizes. To address this issue, Chen et al. [57] propose an Adaptive patch-word Matching (AdaMatch) model which matches texts from the report with adaptive patches in the medical image to provide explainability. It utilizes an Adaptive Patch extraction module to dynamically capture abnormal regions of varying sizes and positions. To further improve its explainability, they propose a cyclic CXR-report generation pipeline (AdaMatch-Cyclic), which can perform both CXR-to-report and report-to-CXR generation based on patch-word alignment.
@@
Additionally, Tanida et al. [58] propose a region-guided radiology report generation model that describes individual regions to form the final report. Specifically, an object detector is used to identify and extract visual features of distinct anatomical regions in the chest X-ray image, and then the abnormal regions are selected and passed to a pre-trained LLM like GPT-2 which generates sentences for each selected region. Finally, the generated sentences are post-processed to remove duplicates and concatenated to form the final report. This process ensures that each sentence in the final report can be linked to an abnormal region in the medical image, which greatly improves explainability.
4-D Medical Q&A
@@
4-D1 Robustness in Medical Q&A
Robustness in medical foundation models for Q&A tasks requires being able to handle errors in the input text prompts as well as be resilient against adversarial and backdoor attacks in medical images. Though some medical foundation models achieve success in their model performance, they exhibit susceptibility to specific text inputs. Sonsbeek et al.  [60] proposes a prefix tuning method for open-ended medical VQA of LLMs for performance improvement, but the experimental results show that the model is heavily sensitive to the prompt structure, as swapping the order of the question embeddings and the visual prefix yields a decrease in performance. To improve the robustness of LLM-based foundation models in medical Q&A, RO-LLaMA [61] incorporates consistency regularization, inspired by NEFTune [80], which adds random noise to the embedding vectors during training. This approach enhances the model’s robustness and generalization abilities when faced with noisy inputs. Additionally, robustness issues are also present in CLIP-based medical Q&A models such as MedCLIP [65] and PubMedCLIP [62], especially if the backbone CLIP-based models themselves have robustness problems. This concern can be inferred by observations in [31], which show that using a modest set of wrongly labeled data and introducing a “Bad-Distance” between the embeddings of clean and poisoned data can lead to successful backdoor attacks.
4-D2 Reliability in Medical Q&A
One significant threat for reliability of medical foundation models is hallucinations, which is hard to evaluate and detect. Although existing works like CHAIR [83] are designed for general foundation hallucination evaluation, they cannot be directly adapted to the medical domain. To address this, the first benchmark and a new evaluation metric named Med-HallMark and MediHall Score respectively, for hallucination detection specific to medical foundation models is proposed to provide baselines for various foundation models in medical image analysis [82]. Meanwhile, a hallucination detector for medical foundation models is proposed for five types of hallucinations: catastrophic, critical, attribute, prompt-induced and minor hallucinations. Similarly, CARES is proposed to comprehensively evaluate the hallucination of Medical Large Vision Language Models (Med-LVLMs) [84]. To ensure reliability, Shaaban et al.  [64] propose MedPromptX and leverage few-shot prompting (FP) to eliminate hallucination in the foundation model for chest X-ray Diagnosis by guiding the output.
Another concern in reliability is truthfulness. A recent study highlights that Medical Visual Question Answering (Med-VQA) systems may be unreliable for medical diagnosis questions and could produce misleading information when existing state-of-the-art foundation models, such as GPT-4, are directly adapted to medical image analysis domains [85]. A good example of this is ChatCAD, which leverages LLMs’ extensive and reliable medical knowledge to provide truthful interactive explanations and advice [63]. However, the truthfulness of ChatCAD is threatened by the restricted scope of applicable imaging domains and the lack of requisite depth in medical expertise during interactive patient consultations. To address these disadvantages of CADs, Zhao et al.  [33] propose ChatCAD+, which incorporates a domain identification module to work with a variety of CAD models. Additionally, outdated medical information can also contribute to incorrect answers. Instead of directly answering medical questions, ChatCAD+ first retrieves knowledge from professional websites to obtain relevant, up-to-date information, which further enhances the truthfulness of its answers.
4-D3 Explainability in Medical Q&A
In high-risk medical environments, the use of black-box foundation models can lead to severe consequences, which may directly impact patient health and safety. Therefore, it is essential to ensure that the decision-making processes of medical foundation models are transparent and explainable. By providing interactive explanations for the advice of the medical image, ChatCAD enhances its explainability for clinical decision-making [63]. With the popular use of GPT-4 in medical fields, providing appropriate prompts that ask the model to explain its answers during medical Q&A can also enhance the explainability [35].
4-D4 Fairness in Medical Q&A
Under-representation of certain groups in the training data or training data imbalance can lead to significantly lower accuracy or effectiveness of the model for certain populations, leading to disparate performance and fairness issues [86]. In medical Q&A systems, these fairness issues can be caused by imbalanced training data in both the medical image and language domains, and can also uniquely arise in the query prompts.
Zhang  et al. [59] propose PMC-VQA, a large-scale medical VQA dataset with 227k VQA pairs of 149k images covering various modalities and diseases. Despite their efforts to construct a comprehensive MedVQA dataset, the authors note the potential presence of biases, which may arise from data collection, annotation (inconsistencies or subjective interpretations from human annotators), or the underlying distribution of the medical images and questions. Figure 3 shows the word clouds for the questions and answers in the training set of the PMC-VQA dataset. The imbalance in word frequencies may be an indication of potential biases in the dataset. For instance, for the word cloud of answers, “X-ray”, “MRI”, and “CT scan” are the most prevalent words, suggesting that models trained on this dataset may perform better on well-represented modalities like X-ray, CT, and MRI, and worse on other imaging modalities.
@@ 4-D4 Fairness in Medical Q&A
Another popular dataset used for medical Q&A tasks is the SLAKE [87] dataset. The distribution of images for different body parts is shown in Figure 4(a). It can be observed that the dataset consists mostly of images of chest, abdomen, and head, with pelvic and neck images only taking up a significantly smaller portion. Thus, the under-representation of pelvic and neck images may lead to worse performance when tested on images from these body regions. Another example is the CheXpert dataset, which is a large public dataset of chest radiographs of 65,240 patients. The age distribution of male and female patients is shown in Figure 4b, which reveals a gender imbalance in the dataset, with a predominance of male patients for most of the age groups, potentially leading to fairness issues in which the model performs better at diagnosing radiographs from male patients.
(a) Distribution of body parts for images in the SLAKE Dataset.
(b) Age Distribution of the CheXpert Dataset.
Figure 4: Data statistics in the medical Q&A datasets.
@@
Chen et al. [12] propose CheXagent, a foundation model designed for chest X-ray interpretation, consisting of three main components: a clinical LLM, a vision encoder, and a vision-language bridger. To evaluate potential fairness concerns present in their proposed model, the authors tested it on a subset of the CheXpert dataset, using chest X-rays from individuals self-reporting as Asian, White, or Black. Resampling with replacement was used to ensure balanced disease prevalence and subgroup representation. The model’s performance was then evaluated for cardiomegaly detection using F1-scores with the prompt “Does this chest X-ray contain cardiomegaly?” Some disparities were observed with the model’s performance on patients of different races and ages. Specifically, F1-scores are highest for the Black subgroup and lowest for the Asian subgroup, reflecting potential inherent differences in the presentation of cardiomegaly across races, and could also be influenced by the limited samples of 14 Black and 30 unique Asian subjects included in the test set. Regarding age, the model performs better for the 65+ age group compared to the 0-65 group, potentially due to a higher prevalence of cardiomegaly in older patients, and age-related physiological differences. To address this fairness issue, one possible way is through curating larger and more diverse datasets, ensuring that models trained using them are more representative and equitable across different patient demographics.
@@
Stereotype bias also leads to fairness concerns, which are typically misleading views and expectations toward particular social groups. The most common stereotype biases include bias towards gender, race, religion, sexual orientation, disability, socio-economic status, and age [86]. Most LLMs are trained on data from the internet, which is riddled with various sources of inherent biases, and LLMs may perpetuate or amplify these existing biases. For example, Bubeck et al. revealed GPT-4’s choice of pronouns amplifies the skewness of the world representation for that occupation [88]. Thus, the integration of LLMs, either directly or after fine-tuning, into medical Q&A systems, requires careful consideration of these inherent bias and fairness issues of the LLMs.
4-E Disease Diagnosis

### v2

@@ 4.1.3 Fairness in Segmentation
Imbalanced training set and variability of clinician needs in segmentation tasks may cause fairness issues. As with base SAM itself, limitations are identified in the ability of MedSAM to perform reliably with underrepresented image modalities. In the context of MedSAM and medical image analysis, the concern comes from the fact that the training set largely consists of CT, MRI, or Endoscopy images. A similar issue is found in the variability of clinician needs in segmentation tasks; not all image modalities should be divided the same way. This presents a concern in the fairness of these segmentation models as they are used in rare clinical scenarios [1].
@@
Performance disparities related to patient BMI and gender are observed in MedSAM. MedSAM [1], a customization of the SAM fine-tuned on over 1M medical image-mask pairs covering 15 image modalities, has become a popular medical image analysis foundation model for segmentation. Xu et al. [39] examines the performance of MedSAM on a kidney CT scan set and looks for disparities in performance depending on patient characteristics. They use Dice similarity scores as their performance metric. No statistically significant relationship between patient age and performance is found, but a statistically significant negative correlation between patient BMI and performance has been noted. Xu et al. [39] also find that images from female patients yield better performance than from male patients.
4.2 Report Generation
@@
4.2.1 Privacy in Report Generation
Deidentication on both medical image data and text reports. Medical applications inherently come with significant concerns regarding patient data privacy. There are standard pipelines to remove sensitive informations from the medical image DICOM files, but how to provent the reports leaking sentitive patient privacy remains lacking standard guidelines. To satisfy privacy requirements, Thawkar et al. [3] use the MIMIC-CXR dataset where the DICOM files of the X-ray data samples are de-identified by the data provider. They further by removing patient information when training their model. To preserve the utility of the model trained on de-identified data, the authors used GPT-3.5-turbo to remove the de-defined symbols “__”, while preserving the original meaning of the reports.
4.2.2 Robustness in Report Generation
@@
Robustness issues on applying to diverse data types and sources. To improve the robustness of these models across different body regions, Zhong et al. [48] propose ChatRadio-Valuer, a model based on LLMs that learns generalizable representations from radiology reports of one institution and adapts to report generation tasks across different body regions (chest, abdomen, muscle-skeleton, head, maxillofacial, and neck). Another work on body-part robustness in report generation proposes a single-for-multiple (S4M) framework to make the model robust to six different body parts (chest, abdomen, knee, hip, wrist, and shoulder) while maintaining its performance or even outperforming other baseline methods [49]. This is achieved by incorporating general radiology knowledge with the radiology-informed knowledge aggregation branch and enhancing the cross-modal alignment by the implicit prior guidance branch. Another robustness issue comes from the heterogeneity of different clinical institutions. To solve this problem, [48] starts with training on radiology reports from one institution to acquire knowledge of particular patterns and representations, and then it undergoes supervised fine-tuning using data from other numerous institutions. This approach ensures that the model can effectively handle a wide range of clinical scenarios and diagnostic tasks, improving its generalization ability and robustness to different institutions in real-world applications.
@@
Robustness to poor data quality. Medical images with blurry boundaries or noise can also lead to robustness concerns. To address the unstable generalization issues, an innovative SAM-guided dual-encoder architecture in MSMedCap is utilized to enable the capture of information with different granularities and overcome the unstable performance due to blurry boundaries, noise, and poor contrast in medical images [50]. Inspired by NEFTune [81], RO-LLaMA also enhances the model’s robustness and generalization abilities when faced with noisy inputs by adding random noise to the embedding vectors during training [62].
4.2.3 Reliability in Report Generation
Inheriting the issue of general foundation models, for example LLMs and VLMs, report generation systems using foundation models also suffer from reliability issues.
Removing prior references from medical reports to improve reliability. Medical foundation models can make hallucinated references to non-existent prior reports when generating medical reports. One reason for this is that these models are trained on datasets of real-world patient reports that inherently refer to prior reports and lead to the incompatibility when solely inputting a medical image [55, 52]. To address this, [55] propose CXR-ReDonE to remove prior references from medical reports by using GPT-3 to rewrite the reports and using token classification to remove words referring to priors.
@@ 4.2.3 Reliability in Report Generation
Improving training data, LLMs, or injecting additional knowledge to improve reliability. In addition, Lee et al. [54] find that their proposed method for chest X-ray report generation may hallucinate reports with nonexistent findings. To mitigate this issue, they suggest some potential solutions, such as using higher quality/quantity training data, larger LLMs, or to strengthen the alignment between image and text modalities. Another work [56] proposes to enhance the reliability of radiology report generation by injecting additional knowledge with the current image in the generation process. Specifically, they fuse the weighted medical concept knowledge and reports for similar images with the current image features. Another study observe that when customizing general-purpose foundation models for medical report generation, they can hallucinate unintended text [51]. They attribute this to the use of cross-entropy loss for language modeling, which leads to the problem of exposure bias. One potential solution to address this issue is to leverage reinforcement learning with human (AI) feedback.
4.2.4 Explainability in Report Generation
@@
Using bounding boxes and LLMs to improve explainability. Explainability is also a crucial concern in medical report generation. To improve the interpretability of report generation systems, Danu et al. [57] propose a two-step approach for generating the findings section of a radiology report from an automated interpretation of chest X-ray images: (1) detecting the abnormalities within the image using bounding-boxes with probabilities, and (2) harnessing the power of LLMs to translate the list of abnormalities into a Findings report. This two-step approach adds interpretability to the framework and aligns it with radiologists’ systematic reasoning during the review of CXRs.
@@
Using adaptive patch-word matching to improve explainability. Using fine-grained vision-language models to provide alignment between image patches and texts which can match parts of the generated report with specific regions in the medical image can also improve explainability. However, using image patches with fixed sizes may result in incomplete representation of lesions which can occur in varying sizes. To address this issue, Chen et al. [58] propose an Adaptive patch-word Matching (AdaMatch) model which matches texts from the report with adaptive patches in the medical image to provide explainability. It utilizes an Adaptive Patch extraction module to dynamically capture abnormal regions of varying sizes and positions. To further improve its explainability, they propose a cyclic CXR-report generation pipeline (AdaMatch-Cyclic), which can perform both CXR-to-report and report-to-CXR generation based on patch-word alignment.
@@
Linking report texts to regions in the medical image to improve explainability. Tanida et al. [59] propose a region-guided radiology report generation model that describes individual regions to form the final report. Specifically, an object detector is used to identify and extract visual features of distinct anatomical regions in the chest X-ray image, and then the abnormal regions are selected and passed to a pre-trained LLM like GPT-2 which generates sentences for each selected region. Finally, the generated sentences are post-processed to remove duplicates and concatenated to form the final report. This process ensures that each sentence in the final report can be linked to an abnormal region in the medical image, which greatly improves explainability.
4.3 Medical Q&A
@@
4.3.1 Robustness in Medical Q&A
Robustness in medical foundation models for Q&A tasks requires being able to handle errors in the input text prompts as well as be resilient against adversarial and backdoor attacks in medical images.
Foundation models can be sensitive to prompt structures. Though some medical foundation models achieve success in the model performance, the susceptibility of these models to specific text inputs still poses a significant challenge. Sonsbeek et al.  [61] proposes a prefix tuning method for open-ended medical VQA of LLMs for performance improvement, but the experimental results show that the model is heavily sensitive to the prompt structure, as swapping the order of the question embeddings and the visual prefix yields a decrease in performance.
Consistency regularization can be used to improve model’s robustness against noisy inputs. To improve the robustness of LLM-based foundation models in medical Q&A, RO-LLaMA [62] incorporates consistency regularization, inspired by NEFTune [81], which adds random noise to the embedding vectors during training. This approach enhances the model’s robustness and generalization abilities when faced with noisy inputs.
CLIP-based models are susceptible to backdoor attacks. Additionally, robustness issues are also present in CLIP-based medical Q&A models such as MedCLIP [66] and PubMedCLIP [63], especially if the backbone CLIP-based models themselves have robustness problems. This concern can be inferred by observations in [31], which show that using a modest set of wrongly labeled data and introducing a “Bad-Distance” between the embeddings of clean and poisoned data can lead to successful backdoor attacks.
4.3.2 Reliability in Medical Q&A
Benchmarks for Hallucinations. One significant threat for reliability of medical foundation models is hallucination, which is hard to evaluate and detect. Although existing works like CHAIR [84] are designed for general foundation hallucination evaluation, they cannot be directly adapted to the medical domain. To address this issue, the first benchmark, Med-HallMark, and a new evaluation metric, MediHall Score, are proposed for hallucination detection specific to medical foundation models, providing baselines for various models in medical image analysis [83]. Meanwhile, a hallucination detector for medical foundation models is proposed for five types of hallucinations: catastrophic, critical, attribute, prompt-induced and minor hallucinations. Similarly, CARES is proposed to comprehensively evaluate the hallucination of Medical Large Vision Language Models (Med-LVLMs) [85]. To ensure reliability, Shaaban et al.  [65] propose MedPromptX and leverage few-shot prompting (FP) to eliminate hallucination in the foundation model for chest X-ray Diagnosis by guiding the output.
Potential issues and solutions of truthfulness. Another concern in reliability is truthfulness. Retrieving knowledge from professional sources can improve the truthfulness of LLMs’ answers. A recent study highlights that Medical Visual Question Answering (Med-VQA) systems may be unreliable for medical diagnosis questions and could produce misleading information when existing state-of-the-art foundation models, such as GPT-4, are directly adapted to medical image analysis domains [86]. An example of this is ChatCAD, which leverages LLMs’ extensive medical knowledge to provide interactive explanations and advice [64]. However, the truthfulness of ChatCAD is threatened by the restricted scope of applicable imaging domains and the lack of requisite depth in medical expertise during interactive patient consultations. To address these issues, Zhao et al.  [33] propose ChatCAD+, which incorporates a domain identification module to work with a variety of CAD models. Additionally, outdated medical information can also contribute to incorrect answers. Instead of directly answering medical questions, Retrieval-Augmented Generated (RAG) is utilized in ChatCAD+ for obtaining relevant, up-to-date information, which further enhances the truthfulness of its answers.
4.3.3 Explainability in Medical Q&A
LLMs can be used to provide explainations. In high-risk medical environments, the use of black-box foundation models can lead to severe consequences, which may directly impact patient health and safety. Therefore, it is essential to ensure that the decision-making processes of medical foundation models are transparent and explainable. By providing interactive explanations for the advice of the medical image, ChatCAD enhances its explainability for clinical decision-making [64]. With the popular use of GPT-4 in medical fields, providing appropriate prompts that ask the model to explain its answers during medical Q&A can also enhance the explainability [35].
4.3.4 Fairness in Medical Q&A
Under-representation of certain groups in the training data or training data imbalance can lead to significantly lower accuracy or effectiveness of the model for certain populations, leading to disparate performance and fairness issues [87]. In medical Q&A systems, these fairness issues can be caused by imbalanced training data in both the medical image and language domains.
Fairness issues can arise from data collection, annotation, or the distribution of the dataset. Zhang  et al. [60] propose PMC-VQA, a large-scale medical VQA dataset with 227k VQA pairs of 149k images covering various modalities and diseases. Despite their efforts to construct a comprehensive MedVQA dataset, the authors note the potential presence of biases, which may arise from data collection, annotation (inconsistencies or subjective interpretations from human annotators), or the underlying distribution of the medical images and questions. Figure 4 shows the word clouds for the questions and answers in the training set of the PMC-VQA dataset. The imbalance in word frequencies may be an indication of potential biases in the dataset. For instance, for the word cloud of answers, “X-ray”, “MRI”, and “CT scan” are the most prevalent words, suggesting that models trained on this dataset may perform better on well-represented modalities like X-ray, CT, and MRI, and worse on other imaging modalities.
@@ 4.3.4 Fairness in Medical Q&A
Fairness issues could arise from imbalanced training data such as SLAKE and CheXpert. Another popular dataset used for medical Q&A tasks is the SLAKE [88] dataset. The distribution of images for different body parts in SLAKE dataset consists mostly of images of chest, abdomen, and head, with pelvic and neck images only taking up a significantly smaller portion. Thus, the under-representation of pelvic and neck images may lead to worse performance when tested on images from these body regions. Another example is the CheXpert dataset, which is a large public dataset of chest radiographs of 65,240 patients. The age distribution of male and female patients reveals a gender imbalance in the dataset, with a predominance of male patients for most of the age groups, potentially leading to fairness issues in which the model performs better at diagnosing radiographs from male patients.
@@
Performance disparities are observed in a foundation model designed for diverse group of datasets. Chen et al. [5] propose CheXagent, a foundation model designed for chest X-ray interpretation, consisting of three main components: a clinical LLM, a vision encoder, and a vision-language bridger. To evaluate potential fairness concerns present in their proposed model, the authors tested it on a subset of the CheXpert dataset, using chest X-rays from individuals self-reporting as Asian, White, or Black. Resampling with replacement was used to ensure balanced disease prevalence and subgroup representation. The model’s performance was then evaluated for cardiomegaly detection using F1-scores with the prompt “Does this chest X-ray contain cardiomegaly?” Some disparities were observed with the model’s performance on patients of different races and ages. Specifically, F1-scores are highest for the Black subgroup and lowest for the Asian subgroup, reflecting potential inherent differences in the presentation of cardiomegaly across races, and could also be influenced by the limited samples of 14 Black and 30 unique Asian subjects included in the test set. Regarding age, the model performs better for the 65+ age group compared to the 0-65 group, potentially due to a higher prevalence of cardiomegaly in older patients, and age-related physiological differences. To address this fairness issue, one possible way is through curating larger and more diverse datasets, ensuring that models trained using them are more representative and equitable across different patient demographics.
@@
Stereotype bias is observed in multi-modal foundation models. Stereotype bias also leads to fairness concerns, which are typically misleading views and expectations toward particular social groups. The most common stereotype biases include bias towards gender, race, religion, sexual orientation, disability, socio-economic status, and age [87]. Most LLMs are trained on data from the internet, which is riddled with various sources of inherent biases, and LLMs may perpetuate or amplify these existing biases. For example, Bubeck et al. revealed GPT-4’s choice of pronouns amplifies the skewness of the world representation for that occupation [89]. Thus, the integration of LLMs, either directly or after fine-tuning, into medical Q&A systems, requires careful consideration of these inherent bias and fairness issues of the LLMs.
4.4 Disease Diagnosis